### v1

# GPT-OSS-20B: A Comprehensive Deployment-Centric Analysis of OpenAI’s Open-Weight Mixture of Experts Model

Abstract
We present a single-GPU (H100, bf16) evaluation of GPT-OSS-20B (Mixture-of-Experts; 20.9B total, $\sim$3.61B active) against dense baselines Qwen3-32B and Yi-34B across multiple dimensions. We measure true time-to-first-token (TTFT), full-decode throughput (TPOT), end-to-end latency percentiles, peak VRAM with past key values (PKV) held, and energy via a consistent nvidia-smi-based sampler.
At a 2,048-token context with 64-token decode, GPT-OSS-20B delivers higher decode throughput and tokens per Joule than dense baselines Qwen3-32B and Yi-34B, while substantially reducing peak VRAM and energy per 1,000 generated tokens; its TTFT is higher due to MoE routing overhead. Concretely, with only 17.3% of parameters active (3.61B of 20.9B), GPT-OSS-20B provides $\sim$31.8% higher decode throughput and $\sim$25.8% lower energy per 1,000 generated tokens than Qwen3-32B at 2,048/64, while using $\sim$31.7% less peak VRAM. Normalized by active parameters, GPT-OSS-20B shows markedly stronger per-active-parameter efficiency (APE), underscoring MoE’s deployment advantages. We do not evaluate accuracy; this is a deployment-focused study. We release code and consolidated results to enable replication and extension.
1 Introduction
The landscape of open-weight large language models has been reshaped by Mixture-of-Experts (MoE) architectures, which activate only a subset of parameters per token and thereby reduce inference cost relative to dense models of similar total size [1]. In this study, we evaluate an open-weight MoE checkpoint, GPT-OSS-20B, which has approximately 20.9B total parameters but only $\sim$3.61B active at inference (17.3% active fraction) [3]. The model was released under a permissive open-source license [2].
While much prior work emphasizes task accuracy, production deployments are constrained by latency, throughput, memory footprint, and energy. This paper addresses those deployment factors with a unified, single-GPU (H100, bf16) evaluation of GPT-OSS-20B against strong dense baselines Qwen3-32B and Yi-34B, reporting true time-to-first-token (TTFT), full-decode throughput (TPOT), end-to-end latency percentiles, peak VRAM with persistent KV, and energy via a consistent nvidia-smi sampler. We further contextualize results using an Active Parameter Efficiency (APE) lens that normalizes performance by the fraction of parameters active at inference. All energy metrics are reported in tokens per Joule and Joules per 1,000 generated tokens for consistency.
1.1 Motivation
The deployment of large language models in production environments requires careful consideration of multiple factors beyond accuracy, including latency, throughput, memory efficiency, and energy consumption. Traditional dense models, while achieving high accuracy, often require significant computational resources that may not be available in resource-constrained environments. Mixture-of-Experts (MoE) architectures offer a promising alternative by activating only a subset of parameters during inference, potentially reducing resource requirements while maintaining competitive performance.
GPT-OSS-20B represents the first open-weight 20.9B parameter MoE model, making it an ideal candidate for comprehensive deployment analysis on a single-GPU (H100, bf16) setup. Understanding its deployment characteristics compared to dense models is crucial for practitioners making architectural decisions in production environments.
1.2 Contributions
Our study makes the following contributions:
  1. 1.
    Unified, reproducible deployment benchmarking. A single-GPU (H100, bf16) harness for latency/throughput (true TTFT, p50/p95, TPOT from full decode), memory (peak VRAM with PKV held), and energy (tokens per Joule, J/1K via stabilized nvidia-smi sampling).
  2. 2.
    Active Parameter Efficiency (APE). A schema-normalized, artifact-robust lens that contextualizes performance by the fraction of parameters active at inference (MoE vs. dense).
  3. 3.
    Memory & KV scaling. Peak-allocator methodology that avoids undercounting transient kernels and reports per-token KV scaling across contexts.
  4. 4.
    Energy efficiency. Apples-to-apples tokens/s, tokens per Joule, and J/1K comparisons under identical decode settings.
  5. 5.
    Ablations. Decoding (greedy vs. sampling), context-length scaling, and precision behavior (bf16 stable for GPT-OSS-20B in our setup); server-stack notes.
  6. 6.
    Safety/governance (qualitative). License and policy overview; we do not claim quantitative safety results.
2 Related Work
2.1 Mixture of Experts Models
Mixture of Experts (MoE) models have emerged as a promising approach to scale language models efficiently. The key insight is that not all parameters need to be active during inference, allowing for larger models with manageable computational requirements. Foundational works include the sparsely-gated MoE layer [1], GShard [9], and the Switch Transformer [10], which demonstrated that MoE can achieve strong performance while significantly reducing active parameter counts.
2.2 Deployment-Centric Evaluation
While most evaluations focus on accuracy metrics, deployment characteristics are crucial for real-world applications. Efficiency aspects such as latency, throughput, and energy consumption have been highlighted in recent surveys of resource-efficient LLMs [4] as well as standardized benchmarks such as MLPerf Inference [5]. However, systematic comparisons of MoE versus dense models under deployment constraints remain limited.
2.3 Open-Weight Model Evaluation
Recent evaluations of open-weight models such as Mistral [6], Qwen [7], and Yi [8] have emphasized accuracy benchmarks. Deployment-centric trade-offs, particularly between MoE and dense models, remain underexplored.
3 Methodology
3.1 Experimental Setup
Hardware. All measurements were taken on a single NVIDIA H100 GPU (bf16), with no sharding or offload. We pin to one device, clear caches between runs, and hold the persistent KV cache (PKV) in memory during measurement.
Software. PyTorch (bf16), transformers, CUDA, and nvidia-smi. We use minimal Python scripts for latency, memory, and energy to avoid framework-induced variability.
Models. We evaluate one open-weight MoE model and two dense baselines:
  • GPT-OSS-20B (Mixture-of-Experts; 20.9B total, $\sim$3.61B active parameters).
  • Qwen3-32B (Dense; 32B total, 32B active).
  • Yi-34B (Dense; 34B total, 34B active).
3.2 Prompting and Context Control
When a tokenizer exposes a chat template, we apply it (apply_chat_template) and then control the post-template context length exactly by trimming/padding to $c$ tokens. If no template exists, we use a simple instruction/completion wrapper and enforce the same post-template $c$. This yields apples-to-apples prefill cost across models.
3.3 Latency Measurement
We report:
  • True TTFT (ms): time to generate one token including prefill. Median of 5 independent trials at the target context $c$.
  • Decode E2E latency: wall-clock time for generating $g$ new tokens, repeated $N$ times per $(c,g)$ pair. We report p50 and p95 over the $N$ runs.
  • TPOT (tok/s): tokens-per-second over the full decode segment, computed from wall time; we report the median across runs. When a TPOT value is suspiciously equal to $1/\text{TTFT}$, we correct it by recomputing from the measured decode wall time.
3.4 Peak Memory Measurement
We measure GPU memory using the CUDA allocator:
  • After a short warm-up (discarded), we reset peak stats and run the actual decode while keeping PKV alive.
  • We read max_memory_allocated() as Peak VRAM, which captures KV cache plus transient kernels.
  • We also compute a peak-based estimate of KV contribution by contrasting peak with pre-decode allocation; this avoids undercounting from “after-run” snapshots.
3.5 Energy Measurement
Energy is sampled with nvidia-smi before/after each decode run and averaged over short, repeated runs:
  • We record instantaneous power (W) around each generation, average across the run, and aggregate over $N$ runs.
  • We report tokens/s, tokens per Joule, and J/1K decoded tokens. The per-1K figure is normalized by decoded tokens ($g$), matching our throughput/latency focus.
  • Power from nvidia-smi is approximate; using identical sampling and run structure across models makes the comparisons reliable.
3.6 Active Parameter Efficiency (APE)
APE contextualizes performance by the fraction of parameters active at inference:
APE-TPOT $=\frac{\text{TPOT}}{\text{Active Params (B)}},$
APE-Energy $=\frac{\text{Tokens/J}}{\text{Active Params (B)}},$
APE-1/TTFT $=\frac{1/\text{TTFT (s)}}{\text{Active Params (B)}},$
TPOT/GB $=\frac{\text{TPOT}}{\text{PeakMem (GB)}}\quad\text{(auxiliary, not an APE term)}.$
3.7 Ablation Protocols
We run controlled ablations under the same harness:
  • Decoding: greedy vs. sampling (top-$p$, top-$k$; temperature sweeps).
  • Context scaling: $c\in\{512,1024,2048,4096\}$ with fixed $g$.
  • Precision: bf16 (primary). FP16/FP32 attempts are reported when supported; bf16 is stable for our MoE runs.
  • Serving stack: direct transformers runs; comparisons with alternative servers (e.g., vLLM) require separate deployment and are not included in the core numbers.
4 Results
Unless otherwise noted, all core results use a context length of 2,0482{,}048 tokens and $64$ decoded tokens, evaluated on a single H100 GPU in bf16. Prompts follow each model’s official chat template, with exact post-template length enforced by trimming or padding. We sweep context lengths $\{128,512,1024,2048\}$ (and $4096$ in ablations), vary decoding settings and precision, and release CSVs and scripts in the repository. Active Parameter Efficiency (APE) is computed with schema normalization and corrected for TPOT artifacts when detected. Minor variances in metrics reflect medians across runs; see CSVs for raw data.
4.1 Latency Analysis
Unless noted, we report single-GPU (H100, bf16), exact post-template contexts, and 64 decoded tokens. TTFT is the time to generate one token (including prefill); TPOT is median tokens/s over the full decode; p50/p95 are end-to-end wall times.
Table 1: Latency at 2048 context and 64 generated tokens. TTFT includes prefill.
Model TTFT (ms) p50 (ms) p95 (ms) TPOT (tok/s)
GPT-OSS-20B $459.72$ $2056.73$ $2060.72$ $31.27$
Qwen3-32B $369.46$ $2738.99$ $2747.82$ $23.73$
Yi-34B $368.34$ $2428.80$ $2434.52$ $26.30$
Table 2: Throughput (TPOT, tok/s) vs. context length (64 generated tokens).
Context GPT-OSS-20B Qwen3-32B Yi-34B
128 $39.79$ $26.56$ $31.66$
512 $38.18$ $25.94$ $30.55$
1024 $36.20$ $25.02$ $28.95$
2048 $31.27$ $23.73$ $26.30$
Table 3: TTFT (ms) vs. context length (64 generated tokens). TTFT includes prefill.
Context GPT-OSS-20B Qwen3-32B Yi-34B
128 $61.02$ $46.09$ $54.43$
512 $188.98$ $111.59$ $110.24$
1024 $203.56$ $193.17$ $192.46$
2048 $459.72$ $369.46$ $368.34$
[…]
Model $\Delta$ TPOT (%) TPOT@128 TPOT@2048
GPT-OSS-20B 21.4-21.4 $39.79$ $31.27$
Qwen3-32B 10.7-10.7 $26.56$ $23.73$
Yi-34B 17.0-17.0 $31.66$ $26.30$
[…]
4.2 Memory Analysis
We measure peak VRAM via the CUDA allocator (max_memory_allocated) while the past-key/value (PKV) cache is held, immediately after the full decode completes. Inputs are post-template trimmed/padded to an exact context length, ensuring identical token counts across models. One short warm-up precedes measurement to stabilize kernel paths; peak stats are then reset and re-measured on the real run.
Table 5: Peak VRAM at context =2,048=2{,}048 and decode =64=64 tokens (allocator peak with PKV alive). Lower is better. $\Delta$ vs. GPT-OSS” is an absolute MB gap; “% less vs. GPT-OSS” uses the baseline’s memory as the denominator.
Model Peak VRAM (MB) $\Delta$ vs. GPT-OSS (MB) % less vs. GPT-OSS
GPT-OSS-20B 43 46143\,461 0 $0.00$
Qwen3-32B 63 65063\,650 20 18920\,189 $31.71$
Yi-34B 66 45966\,459 22 99822\,998 $34.60$
Findings.
At 2,048 context, GPT-OSS-20B uses 31.71% and 34.60% less peak VRAM than Qwen3-32B and Yi-34B, respectively (Table 5), calculated relative to each model’s peak VRAM. In absolute terms, this corresponds to reductions of $\sim$20.2 GB and $\sim$23.0 GB versus the dense 30–34B baselines on a single H100. These savings stem from GPT-OSS-20B’s MoE architecture, which requires less baseline memory (39.9 GB) compared to Qwen3-32B (62.5 GB) and Yi-34B (65.6 GB), despite a larger KV cache footprint, measured with identical post-template token counts.
Caveats.
Allocator peaks include transient workspace from fused kernels and can vary slightly across library versions. We mitigate this by (i) a dedicated warm-up, (ii) resetting peak counters prior to the measured run, and (iii) holding PKV alive through the readout. Results reflect a single-GPU (H100, bf16) configuration with no sharding or offload.
4.3 Energy Analysis
We estimate energy using nvidia-smi power samples on repeated short runs (one warm-up), then compute throughput (TPOT), tokens per watt, and Joules per 1K generated tokens from the measured wall time. We report medians over repeats and enforce the same post-template context as in latency/memory.
Table 6: Energy metrics at context =2,048=2{,}048, decode =64=64. Higher is better for tokens/W; lower is better for J/1K generated tokens.
Model TPOT (tok/s) Tokens/W J/1K (J)
GPT-OSS-20B $31.27$ $0.102$ $9764.2$
Qwen3-32B $23.73$ $0.076$ 13 155.113\,155.1
Yi-34B $26.30$ $0.074$ 13 464.313\,464.3
Findings (ctx=2K).
Relative to Qwen3-32B, GPT-OSS-20B delivers +31.8% higher TPOT, +34.7% higher tokens/W, and -25.8% lower J/1K generated tokens (Table 6). Versus Yi-34B, it shows +18.9% TPOT, +37.9% higher tokens/W, and -27.5% lower J/1K.
Table 7: Energy per 1K generated tokens (J) across post-template context lengths (decode =64=64).
Model 128 512 1024 2048
GPT-OSS-20B 6672.6 7318.5 7976.0 9764.2
Qwen3-32B 10257.0 12295.2 12000.4 13155.1
Yi-34B 10233.6 11460.7 12098.0 13464.3
Context trends.
[…]
Caveats.
(i) nvidia-smi is a coarse, device-level sampler (micro-bursts not captured). (ii) Short-run medians make absolute J/1K approximate; we emphasize relative deltas under identical settings. (iii) All results are single-GPU (H100, bf16), no sharding/offload.
4.4 Active Parameter Efficiency (APE)
APE normalizes performance by the fraction of parameters active at inference:
APE-TPOT $=\frac{\text{TPOT}}{\text{Active Params (B)}},\qquad\text{APE-1/TTFT}=\frac{1/\text{TTFT}}{\text{Active Params (B)}},$
APE-Energy $=\frac{\text{Tokens/W}}{\text{Active Params (B)}},\qquad\text{TPOT/GB}=\frac{\text{TPOT}}{\text{PeakMem (GB)}}.$
Table 8: APE at ctx=2,048=2{,}048, gen=64=64 (per-active-parameter view).
Model Active (B) APE-TPOT APE-1/TTFT APE-Energy TPOT/GB
GPT-OSS-20B $3.61$ $8.664$ $0.602$ $0.0283$ $0.719$
Qwen3-32B $32.00$ $0.742$ $0.085$ 0.002 380.002\,38 $0.373$
Yi-34B $34.00$ $0.774$ $0.080$ 0.002 180.002\,18 $0.396$
[…]
4.5 Ablation Studies (GPT-OSS-20B)
We evaluate decoding choices, context-length effects, numeric precision, and serving stack. Unless noted, we generate 64 new tokens with exact post-template contexts and report median throughput (p50 tok/s) and p50 wall time.
Table 9: Decoding parameters (ctx fixed). Median throughput and wall time; $\Delta$ is relative to Greedy.
Method p50 tok/s p50 time (s) $\Delta$ vs Greedy (%)
Greedy $39.45$ $1.622$ $0.0$
Top-p (0.9) $38.71$ $1.653$ 1.9-1.9
Top-k (50) $38.97$ $1.642$ 1.2-1.2
High Temp $38.73$ $1.652$ 1.8-1.8
Low Temp $38.61$ $1.657$ 2.1-2.1
Takeaway. Sampling reduces throughput by only $\approx$1–2% vs. Greedy with near-identical wall time.
Table 10: Context-length scaling (Greedy). Throughput drops gradually as context grows.
Context p50 tok/s $\Delta$ vs 512 (%)
512 $36.29$ $0.0$
1024 $34.53$ 4.8-4.8
2048 $30.25$ 16.6-16.6
4096 $21.78$ 40.0-40.0
Takeaway. Decode speed degrades smoothly with higher prefill cost; at 4K context it is $\sim$40% below the 512-token baseline.
Table 11: Numeric precision sweep (Greedy). BF16 is stable; FP16/FP32 attempts failed in this harness.
Precision p50 tok/s Notes
BF16 $38.66$ Default
FP16 Failed to run (dtype mismatch)
FP32 Failed to run (dtype mismatch)
Table 12: Serving framework. vLLM requires a separate server (not included here).
Framework p50 tok/s p50 time (s) Notes
Transformers $38.02$ $1.683$ Direct generate()
vLLM Not run (server out of scope)
Overall.
Across decoding strategies, throughput varies by only $\sim$2%. Longer contexts reduce throughput in line with higher prefill cost (table 10). BF16 runs cleanly; other dtypes failed in this harness (table 11). All ablations use exact post-template contexts and identical generation length for fair comparison.
4.6 Safety and Governance (Qualitative)
This section summarizes documentation for each model—license class, presence of a usage-policy link, governance notes, and listed safety features. It is a qualitative, documentation-only review; no quantitative harmlessness/jailbreak testing was run in this study.
Table 13: Safety & governance overview from model cards and metadata. “Policy link” and “Card link” indicate whether a direct URL was present. Safety features are as documented.
Model License (class) Policy link Card link Safety features (as documented)
GPT-OSS-20B Apache 2.0 (Permissive) Yes Yes Designed to follow OpenAI’s safety policies; harmony response format; governance: OpenAI, Safety Advisory Group (SAG)
Qwen3-32B Qwen License (Restricted) No Yes Safety training during development; governance: Alibaba
Yi-34B Apache 2.0 (Permissive) Yes Yes Data compliance checking during training; governance: 01.AI
Key points.
(1) License classes vary; confirm exact terms before deployment. (2) Usage-policy URLs were not always present; model-card links exist for all. (3) Models list qualitative safety features where available; effectiveness not measured here.
Recommendations.
Provide authoritative citations for license/policy/governance, augment with quantitative safety evaluation (harmlessness and jailbreak suites), and document any deployment-time filters or guardrails alongside model results.
5 Conclusion
At 2048-token contexts, GPT-OSS-20B delivers higher throughput, lower energy per 1,000 generated tokens, higher tokens per Joule, and substantially lower peak memory than dense baselines Qwen3-32B and Yi-34B in our single-GPU (H100, bf16) setup, though with higher TTFT due to its MoE architecture. APE complements raw metrics by normalizing for active parameters, highlighting GPT-OSS-20B’s efficiency per active parameter. We release code and consolidated results to facilitate replication.
Reproducibility
  • Single-GPU (H100), bf16; exact post-template context with trim/pad
  • TTFT: 1-token generation including prefill; median-of-5
  • Decode: p50 over 5 runs; TPOT from full decode (median)
  • Memory: peak allocator while PKV alive (max_memory_allocated)
  • Energy: repeated short runs; tokens per Joule and J/1K normalized by generated tokens
  • APE: schema-normalized; TPOT artifact correction when detected
  • Code/results: https://github.com/deepdik/GPT-OSS-20B-analysis